An electron is released from rest at a perpendicular distance of 8.2 cm from a line of charge on a very long nonconducting rod. That charge is uniformly distributed, with 6.0 µC per meter. What is the magnitude of the electron's initial acceleration?
__________________ m/s2
Answer #1

electric field from a infinite line charge is
E = 2kλ/z

where λ is the charge per unit length, and z is the distance to the line

F = qE = ma

a = qE/m = 2kλq/zm = 2*9*10^9*6.0*10^-6*1.6*10^-19/(0.082*9.1*10^-31)

= 3.7*10^17 m/s^2